### v1

# Benchmarking Performance, Explainability, and Evaluation Strategies of Vision-Language Models for Surgery: Challenges and Opportunities

Abstract
Minimally invasive surgery (MIS) presents significant visual and technical challenges, including surgical instrument classification, and understanding surgical action involving instruments, verbs and anatomical targets. While many machine learning-based methods have been developed for surgical understanding, they typically rely on procedure and task-specific models trained on small, manually annotated datasets. In contrast, the recent success of vision-language models (VLMs)—models trained on large volumes of raw image-text pairs—has demonstrated strong adaptability to diverse visual data and a range of downstream tasks. This opens meaningful research questions: how well do these general-purpose VLMs perform in the surgical domain, In this work, we aim to explore those questions by benchmarking several VLMs across diverse surgical datasets, including general laparoscopic procedure and endoscopic submucosal dissection to assess their current capabilities and limitations. Our benchmark work reveals key gaps in the models’ ability to correctly link language to the right areas in the surgical scene in a consistent way.
1 Introduction
Minimally invasive surgery (MIS) offers substantial benefits for patients, including smaller incisions, reduced bleeding and pain, and faster recovery times. However, MIS remains technically challenging for surgeons due to a limited field of view, reduced depth perception, and constrained hand-eye coordination. To address these challenges, many studies have proposed machine learning algorithms for classifying and segmenting surgical instruments and tissues, recognizing procedural phases, and assessing surgical skill Twinanda et al. (2016a); Jin et al. (2020); Zisimopoulos et al. (2018). This approach helps interpret surgical progress, inform potential risks and support real time assistant for surgeons. Despite these efforts, most existing approaches still rely on training task-specific models using relatively small, manually labeled datasets created by clinical experts. These datasets are costly to annotate and limited in scale, which hinders generalization. Moreover, surgical scenes present inherent visual challenges, including varying instrument distances and poses, deformable and visually similar tissues, occlusions, motion blur, and poor lighting conditions.
Minimally invasive surgery (MIS) is routinely performed under video guidance, meaning surgeons rely heavily on visual feeds to carry out procedures. As a result, large volumes of surgical video data are generated daily, much of which already exists but remains underutilized for computational analysis and model training. Additionally, numerous instructional and teaching materials are publicly available online. Identifying the value of raw surgical data remains an open question, particularly in how to effectively leverage them. Recently, the success of vision-language models (VLMs) has demonstrated the potential of learning from large-scale, unstructured data. Models like CLIP Radford et al. (2021) learn visual perception from supervision embedded in natural language, using paired images and text instead of meticulously labeled datasets. This paradigm aligns naturally with the use of raw surgical data and opens promising avenues for surgical AI. When paired with audio, reports or lecture material, large-scale surgical video recording could help overcome the limitations of current models that are restricted to small, manually annotated datasets. Moreover, the revolution in natural language processing enables a more comprehensive understanding of the surgical environment, far beyond simple classification tasks. VLMs have demonstrated remarkable capabilities in processing and reasoning over both visual (e.g., images, videos) and textual (language) data. By combining computer vision and natural language, they allow machines to interpret, reason, and align across modalities Radford et al. (2021); Li et al. (2022); Yuan et al. (2024a, b, 2023).
Figure 1: Examples of VLM doing surgical instrument classification task using CLIP-ViT-B/32 with Grad-CAM to visualize where VLM is looking at. Images (a) and (b) are sampled from the CholecT80-BoundingBox dataset Twinanda et al. (2016b), while (c) and (d) are from the CoPESD dataset Wang et al. (2024). Image pairs (a) and (b) illustrate common spatial reasoning challenges in surgical environments: instruments may appear at varying distances from the endoscope, partially occluded, or positioned at non-standard angles. Image pairs (c) and (d) highlight the effects of poor lighting conditions and motion blurring, where limited illumination leads the model to focus on irrelevant or background regions.
In scenarios involving surgical robots or AI assistants, it is crucial to detect when a model may fail so that control can be safely handed back to the surgeon. Explainability is especially important. Knowing when and how the model fails matters, as the consequences of failure are significantly higher. This requires the model not only to process complex scenes, but also to focus on the most relevant visual cues. Recent work by Wang et al. Wang et al. (2023) highlights a key concern in vision-language models (VLMs): they often focus on irrelevant background regions rather than task-relevant objects, raising concerns about models making correct predictions for the wrong reasons and undermines model’s reliability and interoperability. As shown in Figure 1.
Given that the application of VLMs in surgical environments is still relatively underexplored. With variety of available VLM architecture and strategy for training, and substantial resources required to collect, this leads to a key question that we aim to explore through benchmarking: How effectively do current vision-language models perform on surgical data? To the best of our knowledge only one work by Rau et al. Rau et al. (2025) has benchmarked VLMs across tasks like classification, detection, segmentation, and risk assessment, these evaluations primarily rely on conventional metrics such as accuracy, F1 score, and mIoU. These metrics fail to capture whether the model’s predictions are grounded in clinically meaningful reasoning. To address this gap, we propose a framework centered on examining model attention during prediction, revealing not only what the model predicts but also whether it attends to the correct regions in the image. Specifically, we incorporate class activation mapping (CAM) techniques such as Grad-CAM Selvaraju et al. (2017), which highlight the spatial regions that contribute most to a model’s decision. These heatmaps allow us to directly assess whether the model’s attention is aligned with the relevant surgical content. By visualizing and quantifying the alignment between model attention and ground-truth regions, in this benchmark work, our main contributions are as follows: (1) We introduce a benchmarking framework for evaluating vision-language models (VLMs) in surgical settings, with a focus on the attention of VLMs beyond conventional classification accuracy. (2) We propose multiple new metrics that quantify whether the focus of a model aligns with relevant regions in the surgical data. (3) We demonstrate a consistent gap between correct predictions and correct attention, raising concerns about the reliability of VLMs in safety-critical surgical contexts.
2 Related Work
2.1 VLM
We benchmark pre-trained vision-language models (VLMs) on tasks involving surgical scene understanding, with a focus on their zero-shot interpretability and spatial alignment. Prior work has employed zero-shot approaches to leverage pretrained models for downstream tasks without the need for fine-tuning, thereby significantly reducing computational resource requirements Alayrac et al. (2022). A notable example is CLIP Radford et al. (2021), which learns aligned image-text representations from large-scale web data and has been widely used in zero-shot classification and retrieval tasks. Techniques such as prompt learning and tuning strategies have also been proposed to improve visual-textual embedding alignment in VLMs Liu et al. (2023); Zhou et al. (2022); Chen et al. (2025). In this work, we adopt a benchmarking approach using frozen models, specifically evaluating their zero-shot generalization capability in the surgical domain.
2.2 VLM for surgical applications
There is growing attention toward joint vision-language learning in surgical domains. In reality, only a handful of studies have begun to explore this area learning joint representations from surgical videos and natural language descriptions, often sourced from both narrated and slide-guided surgical lectures, to obtain more generalization VLMs. SurgVLP Yuan et al. (2023) introduced a large-scale surgical pretraining approach using data from surgical lecture videos. It adopts a CLIP-style training paradigm with a contrastive learning objective to align video clips and their corresponding textual descriptions in a shared embedding space. A follow-up model, HecVL Yuan et al. (2024b), enhances this framework by introducing hierarchical supervision, pairing videos with both free-form text and structured phase- and procedure-level annotations, leading to improved phase recognition. PeskaVLP Yuan et al. (2024a) further extends this line of work by enabling alignment between video and text even when they are not temporally synchronized. Other foundation model such as GSViT Schmidgall et al. (2024), which is also trained on surgical video data using a video prediction objective. While GSViT effectively captures temporal dynamics and achieves strong performance in surgical phase recognition. Most prior work applying VLMs to surgical tasks has focused on temporal modeling and phase recognition. Through our benchmark, we aim to systematically evaluate these models from a different perspective—specifically, their spatial and semantic grounding abilities in zero-shot settings. Rather than replacing prior approaches, our work complements them by identifying limitations across spatial, semantic dimensions. This helps inform future efforts to better adapt and align VLMs with the unique demands of surgical applications.
2.3 Visual explanation methods for Deep learning
While deep nerual networks enable superior performance their lack of decomposability into intuitive and understandable components makes them hard to interpret. Consequently, when today’s intelligent systems fail, they fail spectacularly, disgracefully, without warning or explanation, leaving a user staring at an incoherent output. Several methods have been developed to visualize which regions of an image have the most influence on model predictions. Grad-CAM Selvaraju et al. (2017) highlights important regions by computing the gradients of the output with respect to intermediate feature maps of convolutional networks. For transformer-based models, recent work Chefer et al. (2021) extends this idea by computing gradients with respect to attention layers. Grad-CAM++ Chattopadhay et al. (2018) refines the original approach by improving localization in cases with multiple object instances, using a weighted average of higher-order derivatives. Score-CAM Wang et al. (2020), on the other hand, avoids gradient computation altogether by using the model’s output scores to directly weight the activation maps. This can lead to more stable results, though at a higher computational cost. In our work, we adopt the original Grad-CAM method due to its simplicity and ease of implementation, and use it as a core tool in our benchmark.
2.4 Evaluation metrics for Surgical data
Recent benchmark papers on VLM understanding in surgical environments have focused on the general performance of VLMs. Rau et al. (2025) primarily evaluating classification results for tasks such as tool and anatomy recognition, segmentation, and identifying risk or safety-related information. However, conventional metrics like accuracy or IoU often fail to capture whether the model’s predictions are based on clinically meaningful evidence, especially in high-stakes surgical contexts. To our knowledge, the only work that explicitly questions the suitability of conventional metrics in surgical data is Funke et al. (2023), which focuses on the task of surgical phase recognition and highlights that conventional metrics fail to capture temporal structure such as the correct order and duration of phases. For example, a model that predicts all the right phases but in the wrong sequence, or switches phases too frequently. To address this, they propose using metrics like segmental F1 score and edit distance, which evaluate how well the predicted phase segments match the ground truth in terms of both timing and order. These metrics penalize mistakes such as switching phases too often or predicting the correct phases in the wrong sequence, which are better suited to evaluate whether the model predicts the correct sequence and timing of surgical phases. For quantifying spatial semantic grounding, Chen et al. (2022) use the idea of IoU to evaluate gScoreCAM by measuring the spatial overlap between the model’s binarized attention region and the ground-truth object. However, this approach overlooks how strongly different regions contribute to the prediction, which may carry important information in surgical data due to the complexity of the visual scene. Another limitation is that if the model’s prediction is incorrect, its attention is likely to be misaligned as well—making it difficult to judge whether the attention mechanism is meaningful or reliable when compared to ground-truth regions.
3 Proposed Methods
In this section, we first provide a brief introduction to the goal of this benchmark work in Section 3.1. We then describe conventional tasks on tasks this benchmark will evaluate and Class Activation Mapping (CAM) techniques used to interpret VLMs, including methods designed for ResNet-based architectures as well as those adapted for transformer-based models. Finally, we describe the design of our metrics, which aim to quantify the spatial reasoning capabilities of VLMs in surgical environments and provide straightforward visualizations.
3.1 Problem and Task Definition
In this paper, we examine whether VLM predictions in surgical settings are grounded in the correct visual evidence. Formally, this motivates the following question: Given an image $I\in\mathbb{R}^{H\times W\times 3}$ and a set of natural language prompts $\mathcal{T}={T_{1},T_{2},\dots,T_{n}}$, does the model’s attention align with the image regions in $I$ that are relevant to the predicted prompt? Answering this requires going beyond classification accuracy and assessing whether the model’s visual focus meaningfully supports its predictions. To do so, we first introduce visualization-based methods that evaluate this alignment. Our study focuses on three key tasks: (i) instrument classification and grounding and (ii) surgical action classification via triplet descriptions. These tasks evaluate different aspects of multimodal model understanding, including spatial grounding, contextual reasoning, and object/background discrimination in surgical environments.
Instrument Classification Task:
In the instrument classification task, a vision-language model (VLM) selects the most semantically aligned description from a pool of prompts, each referring to a specific surgical instrument class. Conventional evaluation for this task simply measures the proportion of frames where the predicted instrument matches the ground truth label.
Triplet Classification Task:
In Triplet classification task, vision-language model (VLM) selects the most semantically aligned action description from a pool of candidate triplet prompts to evaluate the model’s ability to infer structured surgical actions. Each prompt follows the template “I use a [instrument] to [verb] the [target]”, representing subject-action-object combinations.Conventional evaluation metrics for this task are similar to the instrument classification task and straightforward, which measures how often each individual component is predicted correctly. Let $x\in\{s^{*},v^{*},o^{*}\}$, then the accuracy for component $x$ is computed as the proportion of frames where the predicted $x$ matches the ground truth.
3.2 Proposed CAM-based metrics
Class Activation Mapping (CAM) methods aim to localize the regions in an image that contribute to a model’s decision. These methods work by analyzing the internal activations of a convolutional neural network and their influence on a target output score. We first talk about how Grad-Cam work in VLM that use ResNet.
3.2.1 Grad-CAM for ResNest VLM
For VLM that use residual net Yuan et al. (2023, 2024b, 2024a), the Grad-Cam formula is defined as following. Let $A\in\mathbb{R}^{K\times H\times W}$ denote the output of a selected convolutional layer, where $K$ is the number of channels, and $H\times W$ is the spatial resolution. Let $A_{k}(i,j)$ denote the activation at position $(i,j)$ in channel $k$. The target output score is denoted by $S_{c}\in\mathbb{R}$, which in the case of VLMs is typically a similarity score between image and text. Let $S_{c}\in\mathbb{R}$ denote the output score corresponding to the model’s prediction or the score of interest. In the case of vision-language models, $S_{c}$ is often defined as the cosine similarity between the image embedding $\phi_{\text{img}}(I)$ and the text embedding $\phi_{\text{text}}(T_{c})$ of a selected class $c$:
$S_{c}=\frac{\phi_{\text{img}}(I)^{\top}\phi_{\text{text}}(T_{c})}{\|\phi_{\text{img}}(I)\|\cdot\|\phi_{\text{text}}(T_{c})\|}.$
Finally, computing similarity scores for a set of candidate text prompts $\{T_{1},T_{2},\dots,T_{N}\}$ and select the one with the highest score:
$\hat{c}=\max_{c\in\{1,\dots,N\}}S_{c},$
where $\hat{c}$ is the predicted class, and Let $S_{\hat{c}}$ denotes the scalar target used for backpropagation in CAM methods.
Grad-CAM computes the gradient of the score $S_{c}$ with respect to the feature map $A$, applies global average pooling to obtain channel importance weights:
$\alpha_{k}=\frac{1}{H\cdot W}\sum_{i=1}^{H}\sum_{j=1}^{W}\frac{\partial S_{\hat{c}}}{\partial A_{k}(i,j)},$
The gradient tells how much each feature map channel $A_{k}$ contributes to the model’s prediction $S_{\hat{c}}$. it does this by averaging the gradient of predicted score with respect to all locations $(i,j)$ in that channel and use it act as a weight that reflects the importance of that channel. Based on it, we can then construct the heatmap as:
$L_{\text{Grad-CAM}}(i,j)=\text{ReLU}\left(\sum_{k=1}^{K}\alpha_{k}\cdot A_{k}(i,j)\right).$ (1)
For VLM that use transformer, we sue the following method.
3.2.2 Gradient-Based Attention Rollout for CLIP-ViT
Grad-CAM can be adapted to Vision Transformers used in vision-language models (VLMs), specifically CLIP Radford et al. (2021) with a ViT image encoder. Following the method from Chefer et al. (2021), instead of using convolutional feature maps, we leverage the self-attention matrices and their gradients to identify input image patches that contribute most to the similarity between an image and a given text prompt.
Let $\mathcal{T}^{(l)}\in\mathbb{R}^{N\times N}$ denote the self-attention matrix from the $l$-th transformer block in the ViT encoder, where $N$ is the number of tokens, including the [CLS] token. We extract $\mathcal{T}^{(l)}$ after the softmax operation during the forward pass.
We compute the gradient of the similarity score with respect to the attention map:
$\nabla\mathcal{T}^{(l)}=\frac{\partial S_{\hat{c}}}{\partial\mathcal{T}^{(l)}},$
where $S_{\hat{c}}$ is the predicted similarity score corresponding to the selected text prompt in a zero-shot setting. To capture influence on the similarity score, we compute the gradient-weighted attention map:
$\tilde{\mathcal{T}}^{(l)}=\text{ReLU}\left(\nabla\mathcal{T}^{(l)}\odot\mathcal{T}^{(l)}\right),$
where $\odot$ denotes the elementwise (Hadamard) product. This highlights token-to-token interactions that both receive high attention and significantly impact the similarity score. To propagate relevance through the transformer layers, we recursively update a relevance matrix $\mathbf{R}^{(l)}\in\mathbb{R}^{N\times N}$ as follows:
$\mathbf{R}^{(L)}=I,\quad\mathbf{R}^{(l)}=\mathbf{R}^{(l+1)}+\tilde{\mathcal{T}}^{(l)}\mathbf{R}^{(l+1)},$
starting from the final layer $L$, where $I$ is the $N\times N$ identity matrix. This recursive formulation accumulates relevance flow across all layers. Let each patch token indexed by $k\in\{1,\dots,N-1\}$, after propagating to the input layer, we extract the relevance scores from the [CLS] token to each image patch token:
$r_{k}=\mathbf{R}^{(0)}_{\texttt{CLS},\,k},\quad\text{for }k=1,\dots,N-1,$
where $r_{k}$ denotes the relevance of the $k$-th image patch token to the final similarity score.
We reshape and upsample the vector $[r_{1},\dots,r_{N-1}]\in\mathbb{R}^{N-1}$ into the input image resolution using bilinear interpolation. The final heatmap is defined as:
$L_{\text{GradCAM-ViT}}(i,j)=\text{Upsample}\left(\text{reshape}\left([r_{1},\dots,r_{N-1}]\right)\right)_{i,j},$ (2)
where $(i,j)$ indexes pixel locations in the original image. Based on the CAM style methods we now formally define our metrics as following.
3.3 Instrument classification metrics
Notation.
Let $I$ denote an input image of size $H\times W$, and let $T$ be a corresponding text prompt. The model produces a Grad-CAM heatmap $L(i,j)\in\mathbb{R}^{H\times W}$ using (1) or (2), where each pixel $(i,j)$ reflects the spatial importance of that location for the model’s prediction. We normalize $L$ to the range $[0,1]$, and apply a fixed threshold $\tau\in[0,1]$ to etain information about the "strength" of different regions. We now define the predicted $\tau$-attention region as following:
$\mathcal{A}_{\tau}=\left\{(i,j)\in\{1,\dots,H\}\times\{1,\dots,W\}\mid L(i,j)\geq\tau\right\}.$
Let $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ be the set of ground-truth class labels (e.g., surgical instruments) present in the image. For each class $c_{k}\in\mathcal{C}$, let $\mathcal{G}_{c_{k}}^{(r)}\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ denote the set of pixel coordinates in the $r^{\text{th}}$ annotated region of class $c_{k}$, for $r=1,\dots,N_{k}$. Let $\hat{c}$ be the class predicted by the model. We define the set of all annotated pixels (regardless of class) as:
$\mathcal{G}_{\text{all}}=\bigcup_{k=1}^{K}\bigcup_{r=1}^{N_{k}}\mathcal{G}_{c_{k}}^{(r)}\subseteq\{1,\dots,H\}\times\{1,\dots,W\}.$
We define the set of pixels corresponding to the predicted class $\hat{c}$ as:
$\mathcal{G}_{\hat{c}}=\begin{cases}\displaystyle\bigcup_{\begin{subarray}{c}\\ c_{k}=\hat{c}\end{subarray}}\bigcup_{r=1}^{N_{k}}\mathcal{G}_{c_{k}}^{(r)},&\text{if }\hat{c}\in\mathcal{C}\\ \emptyset,&\text{otherwise}\end{cases}$
(1) Thresholded Attention Coverage ($\tau$-AC):
This metric measures how much of the model’s attention region overlaps with any annotated object, regardless of class correctness:
[…]
It reflects the fraction of model attention that lands inside any annotated region. This is a less strict metric that serve as a soft proxy to access whether the model is at least focusing on some semantically meaningful area.
(2) Thresholded Attention Alignment ($\tau$-AA):
This metric evaluates whether the model’s attention aligns specifically with the annotated regions of the predicted class $\hat{c}$. If $\hat{c}\notin\mathcal{C}$, the score is defined to be zero:
[…]
This ensures that only alignment with correctly predicted classes contributes positively to the score. Beyond conventional classification tasks, understanding the surgical action depicted in an image is also critically important in surgical data.
3.4 Triplet evaluation metrics
(1) Action Reasoning Score (ARS):
In the triplet classification task, the traditional and straightforward way of evaluating performance is to compute the F1 classification score for each class. However, in this paper, beyond assessing classification accuracy, we also challenge the spatial reasoning ability of vision-language models (VLMs). When it comes to localizing the bounding box of an action verb, the problem becomes more complex. As discussed in Twinanda et al. (2022), the localization of the action region is derived using class activation maps (CAMs) or attention mechanisms learned by the model. In particular, when an instrument is actively performing an action, the action is spatially grounded near the location of the instrument tip. Due to the lack of labeled action verb bounding boxes, we propose the Action Reasoning Score (ARS). This score measures whether the model’s attention for a predicted action (verb) overlaps with the bounding box of the predicted instrument (subject). Let a predicted triplet be denoted by $(s,v,o)$, where $s\in\mathcal{S}$ is the subject, $v\in\mathcal{V}$ is the verb (e.g., an action such as cut), and $o\in\mathcal{O}$ is the object (e.g., tissue) and ground-truth triplet is $(s^{*},v^{*},o^{*}).$We compute ARS only when: $s=s^{*}\quad\text{and}\quad v=v^{*}.$ Let the image be defined over a pixel grid of height $H$ and width $W$, with each pixel indexed by coordinates $(i,j)\in\{1,\dots,H\}\times\{1,\dots,W\}$. Let $L_{v}(i,j)\in[0,1]$ denote the value of Grad-CAM attention produced by (1) or (2) at pixel $(i,j)$ for the predicted verb $v$. We will later show how to generate the attention heatmap $L_{v}$ in section 4.1.
For a fixed threshold $\tau\in[0,1]$, we define the $\tau$-attention region for verb as following:
$\mathcal{A}_{\tau^{v}}=\left\{(i,j)\mid L_{v}(i,j)>\tau\right\}.$
Recall that $\mathcal{B}_{s}\subseteq\{1,\dots,H\}\times\{1,\dots,W\}$ denote the set of pixels within the bounding box of the predicted instrument $s$. We define the Action Reasoning Score (ARS) as:
$\text{ARS}(s,v)=\begin{cases}\displaystyle\frac{|\mathcal{A}_{\tau^{v}}\cap\mathcal{B}_{s}|}{|\mathcal{A}_{\tau^{v}}|},&\text{ if }s=s^{*},v=v^{*}\\ 0,&\text{otherwise}\end{cases}$ (5)
Note that this score provides a soft measure of how much the verb-specific attention overlaps with the instrument’s spatial region, offering insight into whether the model is focusing on a plausible area for grounding the action. While there is no strict threshold for overlap, a higher value generally indicates better spatial alignment, whereas a near-zero value suggests the model is likely attending to an irrelevant region. A more detailed implementation is provided in the experimental analysis in Section 4.1.
(2) Triplet Classification:
[…]
Given a predicted triplet $(s,v,o)$ and a ground-truth triplet $(s^{*},v^{*},o^{*})$, we compute the proportion of correctly predicted components using:
$\text{TripletMatch}(s,v,o;s^{*},v^{*},o^{*})=\left(\mathbbm{1}_{s=s^{*}},\;\mathbbm{1}_{v=v^{*}},\;\mathbbm{1}_{o=o^{*}}\right)$ (6)
In this work, we focus on evaluating the action reasoning ability of vision-language models (VLMs). Specifically, we are interested only in the classification of the subject and verb components, denoted by the indicators $\mathbbm{1}_{s=s^{*}}$ and $\mathbbm{1}_{v=v^{*}}$ in the ARS score (5).
4 Experiments Set up
4.1 Datasets and implementation
All experiments use a fixed Grad-CAM threshold of $\tau=0.3$ to generate binary attention maps for grounding evaluation. We first introduce all the datasets we used in this work:
Cholec80 BBox:
We use the Cholec80 BBox dataset Jalal et al. (2023); Alshirbaji et al. (2024), an extension of Cholec80 Twinanda et al. (2016b), which provides bounding box annotations for surgical instruments across 80 cholecystectomy videos. This dataset is used for the instrument classification task, allowing us to evaluate model predictions using spatially grounded metrics such as $\tau$-AC (3) and $\tau$-AA (4). It is also used as a source of spatial supervision when computing attention alignment for triplet classification in CholecT45.
CoPESD:
We also include the CoPESD dataset Wang et al. (2024), the first multimodal dataset for endoscopic submucosal dissection (ESD), which contains over 17,000 images with fine-grained, multi-level instrument annotations. This dataset is used for evaluating instrument classification and enables spatial reasoning analysis via the $\tau$-AC and $\tau$-AA metrics.
CholecT45:
[…]
Prompt Design and Evaluation Strategy:
To align with the prompt-based input format of vision-language models (VLMs), we construct a pool of prompts using template-based formulations. For instrument classification, we use prompts template: “an image showing a {instrument class} in use”, covering all possible instrument classes in Cholec80. For triplet-based action reasoning, we use the template “I use a {instrument} to {verb} the {target}. In the instrument classification task, models are not required to predict all classes present in the image. Instead, we generate the heatmap $L_{\text{GradCAM}}$ and $L_{\text{GradCAM-ViT}}$ and evaluate spatial grounding using two metrics: $\tau$-AC, which assesses whether the model attends to any relevant region, and $\tau$-AA, which evaluates whether attention is correctly aligned with the specific object mentioned in the prompt.
For triplet reasoning, we compute the Action Reasoning Score (ARS) only when both the subject (instrument) and verb predictions are correct. In our implementation, similar to the instrument classification setup, we first generate a pool of triplet prompts using the template “I use a {instrument} to {verb} the {target}, covering all possible instrument, verb, and target classes. We then identify frames where both the predicted instrument and verb match the ground truth. For these correctly classified frames, we reuse the predicted verb to construct a secondary prompt of the form “I am performing {predicted_verb}. Finally, we generate its Grad-CAM heatmap $L_{v}$ and compute the action reasoning scores (5).
Due to all datasets used in this study are not yet fully released, we leverage all available video folders and conduct experiments on all released portions.
4.2 Benchmarked Models
In this work, we evaluate four vision-language models (VLMs): CLIP-ViTRadford et al. (2021), HecVLYuan et al. (2024b), PeskaVLPYuan et al. (2024a), and SurgVLPYuan et al. (2023). CLIP-ViT Radford et al. (2021) is the classic and widely adopted architecture for multi-modal learning, trained on large-scale natural image-text pairs. CLIP has demonstrated strong zero-shot performance across various domains, making it a natural reference point for evaluating domain transfer to surgical imagery.
The other three models—SurgVLPPeskaVLP, and HecVL are trained using surgical video datasets with varying levels of supervision, and architectural adaptations. These models are designed specifically to address the challenges of understanding visual-linguistic information in surgical environments, such as tool usage, procedural context, and anatomy. We aim to assess not only a deeper understanding of the reasoning capabilities of VLMs in surgical classification tasks, but also the limitations and trade-offs between general-purpose and task-specific VLMs within the current surgical domain.
Figure 2: Examples of triplet classification tasks using CLIP-ViT with Grad-CAM visualize the verb heatmap $L_{v}$, where CLIP-ViT successfully classifies all images in Panels (a) and (b) with the correct instrument and verb. Panels (a) illustrate Ideal action reasoning, with attention focused near the instrument and high ARS scores. In contrast, panels (b) show failure cases: despite correct triplet predictions, the attention is misplaced on background tissues, leading to ARS scores of 0.00. These examples highlight that some correct classifications are not based on the correct spatial evidences.
5 Results
5.1 Instrument Classification Performance
We first report the thresholded attention coverage ($\tau$-AC) and thresholded attention alignment ($\tau$-AA) scores of instrument classification tasks, $\tau$-AC quantifies the proportion of model attention that overlaps with any annotated tool region. $\tau$-AA further restricts the evaluation to the tool class predicted by the model.
Model Video-41 Video-42 Video-43 Video-44 Video-45 CoPESD
$\tau$-AC $\tau$-AA $\tau$-AC $\tau$-AA $\tau$-AC $\tau$-AA $\tau$-AC $\tau$-AA $\tau$-AC $\tau$-AA $\tau$-AC $\tau$-AA
CLIP-ViT 0.2022 0.0356 0.2483 0.0893 0.2926 0.0445 0.2038 0.0484 0.2395 0.1021 0.3405 0.0680
HecVL 0.0574 0.0097 0.0953 0.0334 0.0824 0.0199 0.0503 0.0162 0.0565 0.0285 0.1586 0.0653
PeskaVLP 0.0918 0.0333 0.1301 0.0373 0.1252 0.0341 0.1030 0.0306 0.1225 0.0637 0.2145 0.0714
SurgVLP 0.0659 0.0111 0.0917 0.0164 0.1215 0.0275 0.0736 0.0118 0.0645 0.0203 0.2247 0.1522
Table 1: Spatial grounding performance for instrument classification per video folder (video-41 to video-45) in the Cholec8k Bounding Box dataset, and overall on CoPESD at threshold $\tau=0.3$. Each cell contains either Attention Correctness ($\tau$-AC) or Attention Agreement ($\tau$-AA).
Table 1 shows the spatial reasoning performance of five vision-language models across two datasets. Overall, all VLMs tested so far perform poorly under complex surgical environments. CLIP consistently outperforms the other models in both $\tau$-AC and $\tau$-AA metrics across most videos, achieving the highest $\tau$-AC of 0.3405 and $\tau$-AA of 0.0680 on the CoPESD dataset, and a $\tau$-AC of 0.2395 and $\tau$-AA of 0.1021 on the Cholec8k Bounding Box dataset. In contrast, despite being trained on surgical lectures, HecVL, PeskaVLP, and SurgVLP appear to focus more on understanding the overall surgical phase rather than performing fine-grained, frame-level instrument classification.
5.2 Triplet Classification
We compute the Action Reasoning Score (ARS), which evaluates whether the attention map for a predicted action verb overlaps with the correct instrument’s spatial location. This requires datasets that provide both triplet-level annotations and bounding box supervision. We combine the CholecT45 dataset (triplet annotations) and the Cholec80BoundingBox dataset (bounding boxes); only two video folders—video-42 and video-43 are common to both.
Video SurgVLP PeskaVLP HecVL CLIP-ViT
ARS V/T (%) Z/V (%) ARS V/T (%) Z/V (%) ARS V/T (%) Z/V (%) ARS V/T (%) Z/V (%)
Video-42 0.0669 30.5 59.1 0.0744 50.9 43.4 0.0736 68.1 46.9 0.0895 24.2 20.6
Video-43 0.0535 19.7 53.0 0.0673 24.1 51.1 0.0832 24.3 45.9 0.1486 9.5 19.2
[…]
Based on Table 2, CLIP-ViT shows a better ability to pay attention to the right location when reasoning about the action verb. For example, it achieves the highest ARS scores among all models, with 0.0895 on Video-42 and 0.1486 on Video-43. However, it struggles to classify surgical instruments accurately, as shown by its low valid prediction rates: only 24.2% of frames in Video-42 and 9.5% in Video-43 have both the instrument and verb correct. On the other hand, SurgVLP, PeskaVLP, and HecVL show lower ARS scores (e.g., SurgVLP: 0.0669 on Video-42 and 0.0535 on Video-43) but relatively higher classification accuracy, with HecVL achieving valid frame rates of 68.1% and 24.3% on Videos 42 and 43, respectively. However, among these valid predictions, a large portion have zero ARS scores — for instance, 59.1% of valid frames in SurgVLP (Video-42) and 53.0% in Video-43 show no spatial alignment with the action verb.
5.3 Discussion
Based on the results, we observe that the overall performance of vision-language models (VLMs) on surgical data is generally limited. For CLIP, we observed that it can focus on the correct object and achieves relatively good ARS scores for the verb associated with the instrument. This suggests that vision-language models (VLMs) trained on diverse, large-scale datasets can still perform relatively well on downstream tasks, leveraging their generalized understanding from broad visual concepts. However, we also observed a notable gap between classification accuracy and ARS scores, indicating that the models often make correct predictions without actually attending to the correct spatial regions.
Another problem is the discrepancy between classification and attention, for SurgVLP, PeskaVLP, and HecVL. We believe the main issues is that those three models are trained to focus more on the high-level structure of surgical videos. As a result, they may rely heavily on phase-level patterns and struggle to capture fine-grained, frame-level details. These models often learn from broader contextual cues such as surgical summaries or accompanying audio rather than detailed spatial features within individual frames. This reliance on high-level context means they may overlook the specific regions relevant to the text prompt. We provide Figure 2 to visually illustrate this issue and make the spatial misalignment more evident.
6 Conclusion
Our work introduces a benchmarking framework for evaluating vision-language models (VLMs) in surgical settings, with a focus on spatial attention and grounding beyond standard accuracy metrics. We propose new evaluation measures that assess how well model attention aligns with corresponding relevant regions, providing deeper insight into model behavior. Our findings reveal not only the generally low performance of VLMs on surgical data, but also a consistent disconnect between correct predictions and appropriate visual focus. These issues raise critical concerns about the reliability and interpretability of current VLMs in surgical applications. Even models trained specifically on surgical data struggle when faced with frame-level, fine-grained tasks such as action recognition and instrument classification.